OpenStudy (jazzyfa30):

cos theta =(squareroot70)/4 and 0 degrees

OpenStudy (anonymous):

find \[\sin \theta ,then \sin 2\theta=2\sin \theta \cos \theta \]

OpenStudy (jazzyfa30):

@ganeshie8

ganeshie8 (ganeshie8):

use below to find value of \(\sin \theta\) : \(\sin ^2 \theta + \cos^2 \theta = 1\)

ganeshie8 (ganeshie8):

plugin the give \(\cos \theta \) above and solve \(\sin \theta \)

OpenStudy (jazzyfa30):

\[1-\sqrt{70}/4\]

OpenStudy (anonymous):

\[\frac{ \sqrt{70} }{ 4 }>1, which~ is ~\not` correct ~as \left| \cos \theta \right|\le 1\]

OpenStudy (jazzyfa30):

\[1-(\sqrt{70}/4)^2\]

OpenStudy (anonymous):

check your statement.

OpenStudy (jazzyfa30):

im lost^^^

OpenStudy (jazzyfa30):

thanks

ganeshie8 (ganeshie8):

your question has a typo : cos theta =(squareroot7\(\color{Red}{0}\))/4 and 0 degrees<theta<90 degrees

ganeshie8 (ganeshie8):

check the original question again, that \(\color{Red}{0}\) doesnt belong there ^

OpenStudy (jazzyfa30):

im done yea i see that now thanks for pointing out

ganeshie8 (ganeshie8):

np :)